A manhole cover has an area of 4071.5cm². What is the diameter?
Round to three decimal places.

Answer by Alan3354(69443)

A manhole cover has an area of 4071.5cm². What is the diameter?
---------
Area = pi*r^2 = 4071.5
d = 2*r